PROBABILITY DISTRIBUTION

A plane is being examined by a mechanical engineer. The plane can either be in one of the following states: operational or non operational. Say that the engineer knows the probabilities for the operational and non operational states of the plane. What PROBABILITY DISTRIBUTION would you recommend that would analyze the likelihood of the plane to be operational for 6 out of the upcoming 7 days? Please explain your answer...
Answer #1

Probability Distribution to be used:

Binomial Distribution

n = 7

p = Probability for the operational state of the plane

q = 1- p = Probability for the non - operational state of the plane

P(X=6)=\binom{7}{6}\times q^{1}\times p^{6}=7(1-p)p^{6}